If the marginal propensity to save is 0.35, the multiplier is 2.86.

True

False

Consumption spending is $16 million, planned investment spending is $4 million, unplanned investment spending is $2 million, government purchases are $6 million, and net export spending is $1 million. What is aggregate expenditure?

  

$22 million  

$27 million

$26 million

$29 million

If the multiplier is 5, the marginal propensity to consume must be 0.8.

  

True

False
Answer #1

Solution:-

Answer 1:- Answer is True

The multiplier, m = 1 / (1 - MPC)

                            = 1 / MPS

                            = 1/.35

                            = 2.86

Answer 2:- Answer is $27 Million

Aggregate expenditure (AE) = Consumption spending + investment + government spending + net exports

= $16 Million + $4 Million + $6 Million + $1 Million

= $27 Million

Answer 3:- Answer is True

Spending multiplier = 1 / (1- MPC)  

                       5 = 1 / (1 - MPC)

        5(1 - MPC) = 1

           1 - MPC = 1/5

               1 - MPC = 0.20

             MPC = 1 - 0.20

                    = 0.8

Hence, MPC is 0.8